# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v1

@@ Acknowledgements.
This research was partially supported by the China Scholarship Council (CSC) under grant number 202106220041, the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.References
@@
- Abbasiantaeb et al. (2023) Zahra Abbasiantaeb, Chuan Meng, David Rau, Antonis Krasakis, Hossein A Rahmani, and Mohammad Aliannejadi. 2023. LLM-based Retrieval and Generation Pipelines for TREC Interactive Knowledge Assistance Track (iKAT) 2023. In TREC.
@@
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language Models are Few-Shot Learners. In NeurIPS. 1877–1901.
@@
- Craswell et al. (2019) Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, and Ellen M Voorhees. 2019. Overview of the TREC 2019 Deep Learning Track. In TREC.
@@
- Cronen-Townsend et al. (2002) Steve Cronen-Townsend, Yun Zhou, and W Bruce Croft. 2002. Predicting Query Performance. In SIGIR. 299–306.
@@
- Faggioli et al. (2021b) Guglielmo Faggioli, Oleg Zendel, J Shane Culpepper, Nicola Ferro, and Falk Scholer. 2021b. An Enhanced Evaluation Framework for Query Performance Prediction. In ECIR. Springer, 115–129.
- Faggioli et al. (2022) Guglielmo Faggioli, Oleg Zendel, J Shane Culpepper, Nicola Ferro, and Falk Scholer. 2022. sMARE: A New Paradigm to Evaluate and Understand Query Performance Prediction Methods. Information Retrieval Journal 25, 2 (2022), 94–122.
@@
- Hashemi et al. (2019) Helia Hashemi, Hamed Zamani, and W Bruce Croft. 2019. Performance Prediction for Non-factoid Question Answering. In ICTIR. 55–58.
@@
- Kingma and Ba (2015) Diederik P Kingma and Jimmy Ba. 2015. Adam: A Method for Stochastic Optimization. In ICLR.
@@
- Lavrenko and Croft (2001) Victor Lavrenko and W Bruce Croft. 2001. Relevance-Based Language Models. In SIGIR. 120–127.
@@
- Lu et al. (2016) Xiaolu Lu, Alistair Moffat, and J Shane Culpepper. 2016. The Effect of Pooling and Evaluation Depth on IR Metrics. Information Retrieval Journal 19, 4 (2016), 416–445.
@@
- Meng et al. (2023b) Chuan Meng, Negar Arabzadeh, Mohammad Aliannejadi, and Maarten de Rijke. 2023b. Query Performance Prediction: From Ad-hoc to Conversational Search. In SIGIR. 2583–2593.
- Meng et al. (2024) Chuan Meng, Negar Arabzadeh, Arian Askari, Mohammad Aliannejadi, and Maarten de Rijke. 2024. Ranked List Truncation: From Retrieval to Re-ranking. In SIGIR.
- Meng et al. (2020a) Chuan Meng, Pengjie Ren, Zhumin Chen, Christof Monz, Jun Ma, and Maarten de Rijke. 2020a. RefNet: A Reference-aware Network for Background Based Conversation. In AAAI.
- Meng et al. (2021) Chuan Meng, Pengjie Ren, Zhumin Chen, Zhaochun Ren, Tengxiao Xi, and Maarten de Rijke. 2021. Initiative-Aware Self-Supervised Learning for Knowledge-Grounded Conversations. In SIGIR. 522–532.
- Meng et al. (2020b) Chuan Meng, Pengjie Ren, Zhumin Chen, Weiwei Sun, Zhaochun Ren, Zhaopeng Tu, and Maarten de Rijke. 2020b. DukeNet: A Dual Knowledge Interaction Network for Knowledge-Grounded Conversation. In SIGIR. 1151–1160.
@@
- Sun et al. (2023a) Jiuding Sun, Chantal Shaib, and Byron C Wallace. 2023a. Evaluating the Zero-shot Robustness of Instruction-tuned Language Models. arXiv preprint arXiv:2306.11270 (2023).
@@
- Tunstall et al. (2023) Lewis Tunstall, Edward Beeching, Nathan Lambert, Nazneen Rajani, Kashif Rasul, Younes Belkada, Shengyi Huang, Leandro von Werra, Clémentine Fourrier, Nathan Habib, et al. 2023. Zephyr: Direct Distillation of LM Alignment. arXiv preprint arXiv:2310.16944 (2023).
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought Prompting Elicits Reasoning in Large Language Models. NeurIPS 35 (2022), 24824–24837.
@@
- Zamani et al. (2018) Hamed Zamani, W Bruce Croft, and J Shane Culpepper. 2018. Neural Query Performance Prediction Using Weak Supervision from Multiple Signals. In SIGIR. 105–114.
- Zendel et al. (2023) Oleg Zendel, Binsheng Liu, J Shane Culpepper, and Falk Scholer. 2023. Entropy-Based Query Performance Prediction for Neural Information Retrieval Systems. In QPP++2023. 37–44.
- Zhang et al. (2023b) Xinyu Zhang, Sebastian Hofstätter, Patrick Lewis, Raphael Tang, and Jimmy Lin. 2023b. Rank-without-GPT: Building GPT-Independent Listwise Rerankers on Open-Source Large Language Models. arXiv preprint arXiv:2312.02969 (2023).
@@
- Zhou and Croft (2006) Yun Zhou and W Bruce Croft. 2006. Ranking Robustness: A Novel Framework to Predict Query Performance. In CIKM. 567–574.
- Zhou and Croft (2007) Yun Zhou and W Bruce Croft. 2007. Query Performance Prediction in Web Search Environments. In SIGIR. 543–550.

### v2

@@ Acknowledgements.
This research was partially supported by the China Scholarship Council (CSC) under grant number 202106220041, the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), DPG Media, RTL, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.References
@@
- Abbasiantaeb et al. (2024) Zahra Abbasiantaeb, Chuan Meng, Leif Azzopardi, and Mohammad Aliannejadi. 2024. Can We Use Large Language Models to Fill Relevance Judgment Holes? arXiv preprint arXiv:2405.05600 (2024).
- Abbasiantaeb et al. (2023) Zahra Abbasiantaeb, Chuan Meng, David Rau, Antonis Krasakis, Hossein A. Rahmani, and Mohammad Aliannejadi. 2023. LLM-based Retrieval and Generation Pipelines for TREC Interactive Knowledge Assistance Track (iKAT) 2023. In TREC.
- AI@Meta (2024) AI@Meta. 2024. Llama 3 Model Card. (2024). https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md
@@
- Askari et al. (2023) Arian Askari, Mohammad Aliannejadi, Chuan Meng, Evangelos Kanoulas, and Suzan Verberne. 2023. Expand, Highlight, Generate: RL-driven Document Generation for Passage Reranking. In EMNLP. 10087–10099.
@@
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D. Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language Models are Few-Shot Learners. In NeurIPS. 1877–1901.
@@
- Craswell et al. (2019) Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, and Ellen M. Voorhees. 2019. Overview of the TREC 2019 Deep Learning Track. In TREC.
@@
- Cronen-Townsend et al. (2002) Steve Cronen-Townsend, Yun Zhou, and W. Bruce Croft. 2002. Predicting Query Performance. In SIGIR. 299–306.
@@
- Faggioli et al. (2021b) Guglielmo Faggioli, Oleg Zendel, J. Shane Culpepper, Nicola Ferro, and Falk Scholer. 2021b. An Enhanced Evaluation Framework for Query Performance Prediction. In ECIR. Springer, 115–129.
- Faggioli et al. (2022) Guglielmo Faggioli, Oleg Zendel, J. Shane Culpepper, Nicola Ferro, and Falk Scholer. 2022. sMARE: A New Paradigm to Evaluate and Understand Query Performance Prediction Methods. Information Retrieval Journal 25, 2 (2022), 94–122.
@@
- Hashemi et al. (2019) Helia Hashemi, Hamed Zamani, and W. Bruce Croft. 2019. Performance Prediction for Non-factoid Question Answering. In ICTIR. 55–58.
@@
- Jiang et al. (2023) Albert Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, et al. 2023. Mistral 7B. arXiv preprint arXiv:2310.06825 (2023).
@@
- Khramtsova et al. (2024) Ekaterina Khramtsova, Shengyao Zhuang, Mahsa Baktashmotlagh, and Guido Zuccon. 2024. Leveraging LLMs for Unsupervised Dense Retriever Ranking. arXiv preprint arXiv:2402.04853 (2024).
- Kingma and Ba (2015) Diederik P. Kingma and Jimmy Ba. 2015. Adam: A Method for Stochastic Optimization. In ICLR.
@@
- Lassance and Clinchant (2023) Carlos Lassance and Stéphane Clinchant. 2023. The Tale of Two MSMARCO - and Their Unfair Comparisons. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval. 2431–2435.
- Lavrenko and Croft (2001) Victor Lavrenko and W. Bruce Croft. 2001. Relevance-Based Language Models. In SIGIR. 120–127.
- Lin et al. (2021) Jimmy Lin, Xueguang Ma, Sheng-Chieh Lin, Jheng-Hong Yang, Ronak Pradeep, and Rodrigo Nogueira. 2021. Pyserini: A Python toolkit for reproducible information retrieval research with sparse and dense representations. In SIGIR. 2356–2362.
@@
- Lu et al. (2016) Xiaolu Lu, Alistair Moffat, and J. Shane Culpepper. 2016. The Effect of Pooling and Evaluation Depth on IR Metrics. Information Retrieval Journal 19, 4 (2016), 416–445.
@@
- Ma et al. (2024) Shengjie Ma, Chong Chen, Qi Chu, and Jiaxin Mao. 2024. Leveraging Large Language Models for Relevance Judgments in Legal Case Retrieval. arXiv preprint arXiv:2403.18405 (2024).
@@
- Ma et al. (2021) Yixiao Ma, Yunqiu Shao, Yueyue Wu, Yiqun Liu, Ruizhe Zhang, Min Zhang, and Shaoping Ma. 2021. LeCaRD: A Legal Case Retrieval Dataset for Chinese Law System. In SIGIR. 2342–2348.
@@
- Meng (2024) Chuan Meng. 2024. Query Performance Prediction for Conversational Search and Beyond. In SIGIR.
@@
- Meng et al. (2023b) Chuan Meng, Mohammad Aliannejadi, and Maarten de Rijke. 2023b. System Initiative Prediction for Multi-turn Conversational Information Seeking. In CIKM. 1807–1817.
- Meng et al. (2023c) Chuan Meng, Negar Arabzadeh, Mohammad Aliannejadi, and Maarten de Rijke. 2023c. Query Performance Prediction: From Ad-hoc to Conversational Search. In SIGIR. 2583–2593.
- Meng et al. (2024) Chuan Meng, Negar Arabzadeh, Arian Askari, Mohammad Aliannejadi, and Maarten de Rijke. 2024. Ranked List Truncation for Large Language Model-based Re-Ranking. In SIGIR.
- Mizzaro et al. (2018) Stefano Mizzaro, Josiane Mothe, Kevin Roitero, and Md Zia Ullah. 2018. Query Performance Prediction and Effectiveness Evaluation Without Relevance Judgments: Two Sides of the Same Coin. In SIGIR. 1233–1236.
@@
- Rahmani et al. (2024) Hossein A. Rahmani, Nick Craswell, Emine Yilmaz, Bhaskar Mitra, and Daniel Campos. 2024. Synthetic Test Collections for Retrieval Evaluation. arXiv preprint arXiv:2405.07767 (2024).
@@
- Salemi and Zamani (2024) Alireza Salemi and Hamed Zamani. 2024. Evaluating Retrieval Quality in Retrieval-Augmented Generation. arXiv preprint arXiv:2404.13781 (2024).
@@
- Sun et al. (2023a) Jiuding Sun, Chantal Shaib, and Byron C. Wallace. 2023a. Evaluating the Zero-shot Robustness of Instruction-tuned Language Models. arXiv preprint arXiv:2306.11270 (2023).
@@
- Tay et al. (2022) Yi Tay, Mostafa Dehghani, Vinh Q Tran, Xavier Garcia, Jason Wei, Xuezhi Wang, Hyung Won Chung, Dara Bahri, Tal Schuster, Steven Zheng, et al. 2022. Ul2: Unifying language learning paradigms. In ICLR.
@@
- Upadhyay et al. (2024a) Shivani Upadhyay, Ehsan Kamalloo, and Jimmy Lin. 2024a. LLMs Can Patch Up Missing Relevance Judgments in Evaluation. arXiv preprint arXiv:2405.04727 (2024).
- Upadhyay et al. (2024b) Shivani Upadhyay, Ronak Pradeep, Nandan Thakur, Nick Craswell, and Jimmy Lin. 2024b. UMBRELA: UMbrela is the (Open-Source Reproduction of the) Bing RELevance Assessor. arXiv:2406.06519 [cs.IR]
- Vlachou and Macdonald (2023) Maria Vlachou and Craig Macdonald. 2023. On Coherence-based Predictors for Dense Query Performance Prediction. arXiv preprint arXiv:2310.11405 (2023).
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V. Le, Denny Zhou, et al. 2022. Chain-of-thought Prompting Elicits Reasoning in Large Language Models. NeurIPS 35 (2022), 24824–24837.
@@
- Yan et al. (2024) Le Yan, Zhen Qin, Honglei Zhuang, Rolf Jagerman, Xuanhui Wang, Michael Bendersky, and Harrie Oosterhuis. 2024. Consolidating Ranking and Relevance Predictions of Large Language Models through Post-Processing. arXiv preprint arXiv:2404.11791 (2024).
- Zamani et al. (2018) Hamed Zamani, W. Bruce Croft, and J. Shane Culpepper. 2018. Neural Query Performance Prediction Using Weak Supervision from Multiple Signals. In SIGIR. 105–114.
- Zendel et al. (2023) Oleg Zendel, Binsheng Liu, J. Shane Culpepper, and Falk Scholer. 2023. Entropy-Based Query Performance Prediction for Neural Information Retrieval Systems. In QPP++2023. 37–44.
- Zhang et al. (2024) Hengran Zhang, Ruqing Zhang, Jiafeng Guo, Maarten de Rijke, Yixing Fan, and Xueqi Cheng. 2024. Are Large Language Models Good at Utility Judgments? arXiv preprint arXiv:2403.19216 (2024).
- Zhang et al. (2023b) Shengyu Zhang, Linfeng Dong, Xiaoya Li, Sen Zhang, Xiaofei Sun, Shuhe Wang, Jiwei Li, Runyi Hu, Tianwei Zhang, Fei Wu, et al. 2023b. Instruction Tuning for Large Language Models: A Survey. arXiv preprint arXiv:2308.10792 (2023).
- Zhang et al. (2023c) Xinyu Zhang, Sebastian Hofstätter, Patrick Lewis, Raphael Tang, and Jimmy Lin. 2023c. Rank-without-GPT: Building GPT-Independent Listwise Rerankers on Open-Source Large Language Models. arXiv preprint arXiv:2312.02969 (2023).
@@
- Zheng et al. (2024) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al. 2024. Judging LLM-as-a-Judge with MT-Bench and Chatbot Arena. NeurIPS 36 (2024).
- Zhou and Croft (2006) Yun Zhou and W. Bruce Croft. 2006. Ranking Robustness: A Novel Framework to Predict Query Performance. In CIKM. 567–574.
- Zhou and Croft (2007) Yun Zhou and W. Bruce Croft. 2007. Query Performance Prediction in Web Search Environments. In SIGIR. 543–550.